ADNAN MUHAMMED

Grade 12,

The distance travelled by a body in the nth second is given by the expression 2+3X find the initial velocity and acceleration ?

The distance travelled by a body in the nth second is given by the expression 2+3X find the initial velocity and acceleration ?

Grade:9

1 Answers

poonam
6 years ago
As the formula for distance travelled by the body in nth second is Sn= u+a/2(2n-1)So by comparing it with the given vale i.e. 2+3x. We get u=2 and a/2= 3 so a= 6